Aristotle’s De Interpretatione Chapter 14

Manuel A. Correia

In this paper I examine the authenticity of De Interpretatione Chapter 14, which was brought into question by Porphyry (s. III AD), and still remains without a definitive answer. I give some doctrinal considerations based on recent studies of the logical contents of this chapter. I suggest that it was written by Aristotle but not as a part of De Interpretatione. The Chapter is connected, I argue, with the very point of this treatise, namely, a theory of contradiction, but it corresponds to an earlier development on this matter. My proposal is that (at least the semantic section of) De Int. corresponds to a further answer of the question of Chapter 14, 23a27-31, which actually was made in an earlier state and not as part of the treatise. My main argument is that Aristotle in Chapter 14 disregards his definition of contradiction given in the previous Chapter 6 of this treatise (”I speak of statements as opposite when they affirm and deny the same thing of the same thing”: 17a35-37). It was easy to him, to apply this definition for solving the problem that he puts forward (Ch. 14, 23a27-31), namely: whether the affirmation of the contrary (’this is bad’) or the simple negation (’this is not good’) is the contrary statement of the simple affirmation (’this is good’). Indeed, a simple reflection on what is said in the definition suffices to answer that question: it is only the simple negation the one that denies the same thing of the same thing. The affirmation of the contrary does not deny the same thing of the same thing, but affirms something different (that it is bad, whose negation is ’this is not bad’). Furthermore, (i) Aristotle solves the problem of Chapter 14 by means of a resource that he does not use anymore, neither at An Pr I, 46 nor a bit earlier (De Int. 10, 19b. 19-29), where the relation between these three kinds of proposition is raised again. He argues that a proposition can be more contrary than another. (ii) He also at De Int. 14 is unaware of the exception introduced by singular propositions at De Int. 10, 20a. 23-30, where ’Socrates is just’ is said to be contradictorily denied by ’Socrates is not just’, and by ’Socrates is not-just’ too. Finally, (iii) he at De Int. 14 is unaware of what he says about universal propositions earlier at De Int. 7, 17b.16-25; De Int. 10, 20a23-30, and An Pr I, 46, namely, ’Every man is not-just’ is not the negation of ’Every man is just’, but its contrary. I conclude, then, that the specific relation among the texts that I consider shows that Aristotle is likely the author of De Int. 14. However, since he at De Int. 14 makes a point of something already treated, and implicitly solved, that chapter does not seem to be a genuine part of De Int. Rather, it seems to be earlier than An Pr I, 46 and De Int. 1-13.

Address: Pontificia Universidad Católica de Chile, Instituto de Filosofía, Casilla 316, Santiago 22, CHILE

E-mail: mcorreia@puc.cl